Three identical coils, each of resistance 10ohm and inductance 42mh are connected in star to a 415v, 50hz, 3 phase supply determine a-phase impedance, b-line voltage, c-phase voltage, d-phase current, power factor, e-power dissipated
1
Expert's answer
2020-11-11T08:56:37-0500

Given are following details,

Resistance /phase"(R) =10\varOmega."

Inductance /phase"(L) = 42mH"

Hence Inductive reactance"(X_L) =L\omega=42\times10^{-3}\times2\pi\times50=13.19\varOmega."

a) The perphase impedance"(Z) =R+jX_L=10+j13.19\space\varOmega"


b) Line voltage"(V_L) =415V" (given in question)


c) Phase voltage "(V_P) =\frac{V_L}{\sqrt3}=\frac{415}{\sqrt3}=239.6 V"


d) Phase current"(I_P) =\frac{V_P}{Z}=\frac{239.6}{10+j13.19}=14.17\angle-52.83. A"


Powerfactor "=\cos(52.83) =0.604"


e) Overall power dissipated in the 3phase system "=3\times I_P^2 \times R=3\times14.47^2\times10=6281.43W"